Use the t-distribution to find a confidence interval for a difference in means μ 1 - μ 2 given the relevant sample results. Give the best estimate for μ 1 - μ 2, the margin of error, and the confidence interval. Assume the results come from random samples from populations that are approximately normally distributed.

A 90% confidence interval for μ 1 - μ 2 using the sample results x ¯ 1 = 5.5 , s 1 = 2.3 , n 1 = 10 and x ¯ 2 = 4.5 , s 2 = 2.4 , n 2 = 8

Enter the exact answer for the best estimate and round your answers for the margin of error and the confidence interval to two decimal places.
Answer #1


nalo 525-5 De-4.5 S,=2-3 n28 32=2.4 af=16 At qo% confidence level tag=1.746 The margin of errog (E) = tay. The margin of exto